OpenStudy (anonymous):

How about this one.. 6H=H-20=44-H; H=

OpenStudy (anonymous):

How about it?

jhonyy9 (jhonyy9):

6H = H-20=44-H 5H = -20 -44 = -H 4H = - 64 H= - 64/4 H = - 16